OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 16 raised to the power of 1 over 4 is called the fourth root of 16. (5 points) Part B: The length of a rectangle is 5 units and its width is square root of 5 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well in index from 16 = 2^4 and is that is raised to the 1/4 power \[(2^4)^\frac{1}{4} = 2^?\] what happens..? use the index law for power of a power

OpenStudy (anonymous):

I'm Sorry But What?

OpenStudy (anonymous):

@aaronq

OpenStudy (campbell_st):

well start by rewriting 16... with a base of 2, what is the power...?

OpenStudy (anonymous):

2.....

OpenStudy (campbell_st):

no 2 is the base... so \[2^? = 16\] what value does ? have

OpenStudy (anonymous):

1

OpenStudy (anonymous):

or is it 8

OpenStudy (campbell_st):

no \[2^1 = 2~~~and~~~~2^8 = 256\] you are using powers so what power might you use...

OpenStudy (anonymous):

i do not know .

OpenStudy (anonymous):

4

OpenStudy (campbell_st):

great so \[2^4 = 16\] so raising 16 to the power of 1/4 means \[(16)^\frac{1}{4}\] is you rewrite 16... using 2^4 the problem becomes \[(2^4)^{\frac{1}{4}}\] now when working with powers fraction powers mean radicals so \[16^{\frac{1}{4}} = \sqrt[4]{16}\] and there is an index law that says when finding the power of a power... multiply the powers so rather than use radicals \[(16)^{\frac{1}{4}} = (2^4)^{\frac{1}{4}} = 2^{4 \times \frac{1}{4}}\] so if you multiply the powers what do you get...?

OpenStudy (anonymous):

the problem up top? right

OpenStudy (campbell_st):

no just multiply the powers in \[2^{4 \times \frac{1}{4}}\]

OpenStudy (anonymous):

0.125

OpenStudy (campbell_st):

4 X 1/4 = 1 So it simplifies to 2^1 =2 and when you pick up a calculator \[\sqrt[4]{16} = 2\]

OpenStudy (campbell_st):

for the 2nd question Area = length x width so is \[A = 5 \times \sqrt{5} = 5 \sqrt {5}\] a ration number of an irrational number

OpenStudy (anonymous):

irrational

OpenStudy (campbell_st):

and you have the justification in the calculation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 2 more question :( @campbell_st

OpenStudy (campbell_st):

post them as new questions... and someone will help

OpenStudy (anonymous):

ok